OpenStudy (anonymous):

Factor this expression completely. 36y2 - 1 (6y - 1)2 (6y + 1)(6y - 1) (4y + 1)(9y - 1) (4y - 1)(9y + 1)

OpenStudy (anonymous):

Formula = a^2- b^2 = (a+b)(a-b) So, 36y2 - 1= (6y + 1)(6y - 1)

OpenStudy (anonymous):

i got it

OpenStudy (mathstudent55):

This is called the difference of two squares.